### v2

# Characterization of coherent errors in gate layers with robustness to Pauli noise

Abstract
Characterization of quantum devices generates insights into their sources of disturbances. State-of-the-art characterization protocols often focus on incoherent noise and eliminate coherent errors when using Pauli or Clifford twirling techniques. This approach biases the structure of the effective noise and adds a circuit and sampling overhead. We motivate the extension of an incoherent local Pauli noise model to coherent errors and present a practical characterization protocol for an arbitrary gate layer. Notably, the coherent noise estimation is robust to Pauli noise. We demonstrate our protocol on a superconducting hardware platform and identify the leading coherent errors. To verify the characterized noise structure, we mitigate its coherent and incoherent components using a gate-level coherent noise mitigation scheme in conjunction with probabilistic error cancellation. The proposed characterization procedure opens up possibilities for device calibration, hardware development, and improvement of error mitigation and correction techniques.
I Introduction
Current quantum computers are highly affected by noise [1]. The errors originate from interactions of the devices with their environment [2], unwanted dynamics between the qubits [3], or imperfect control signals [4]. Characterization aims to find error sources and quantify their effect on the performance of a computing processor [5]. The information obtained from the characterization of a device facilitates the development of a less error-prone hardware platform [6], the implementation of schemes for correcting specific errors [7], and the application of noise mitigation protocols [8]. Therefore, noise characterization stands at the core of future technological progress.
Demarcating noise along the line of purity conservation leads to the distinction between coherent and incoherent noise [9]. Coherent noise occurs due to systematic, reversible perturbations such as imperfect calibrations, imprecise control signals, or couplings to low-frequency external fields [10, 11, 12, 13]. Conversely, incoherent noise, e.g., bit-flip errors, is stochastic and typically arises from insufficient isolation of the system from its environment [12], causing a non-reversible loss of information. Since controllability and isolation of a system are generally opposing goals in hardware design [4, 14], distinguishing the structure of the limiting noise is critical to advance quantum computing platforms.
Numerous characterization techniques have been developed to identify and quantify coherent and incoherent noise in quantum processes. Full process tomography [15, 16] and gate-set tomography [17, 18] are common techniques for obtaining a general representation of a physical operation. Although they provide a full representation, they lack in discriminating the noise processes from the actual operation. Furthermore, as they involve the reconstruction of full-system density matrices, without further assumptions, they are experimentally intractable on devices with more than a few qubits due to exponential scaling of the number of operations with the size of the system’s Hilbert space and the resource-intensive post-processing [19]. While some protocols reach scalable noise characterization using principles of randomized benchmarking [20, 21, 22], they are often limited to a few heuristic metrics that do not provide enough insights for existing noise calibration, mitigation, or correction techniques [23]. Characterization approaches that exceed this limitation while still being scalable exist for incoherent noise [24, 7, 25]. These noise reconstruction protocols [26] introduce a model and estimate its parameters by fitting it to the real process. They aim to quantify noise of a gate layer, where a restricted noise correlation length is the central structural assumption. However, those models do not portray the actual noise process completely, as the protocols require the application of probabilistic twirling schemes [27, 28, 20, 29, 30, 31] that convert coherent errors into incoherent ones. Currently, no similar approach to the noise reconstruction protocols exists that is sensitive to coherent errors.
In this article, we introduce a noise model combining coherent and incoherent components. We devise a protocol to characterize the coherent part of the model as a series of unitary transformations, interpretable as qubit rotations. The estimation is independent of the introduced incoherent model components, hardware-agnostic, and scalable to larger systems, assuming a low noise-correlation length. To demonstrate our method, we apply it to characterize a gate layer on a 7-qubit superconducting circuit device provided by IBM Quantum (IBMq) [32]. This allows us to determine the most significant coherent single- and two-qubit errors and their behavior over time. Leveraging these estimation results, we mitigate the coherent errors through a gate-level approach and combine it with probabilistic error cancellation (PEC) [8, 33]. Thereby verifying that our model captures the dominating noise terms. Our work paves the way for advancements in hardware platform development, quantum error correction and mitigation schemes, and the design of noise-aware algorithms.
II Model
An ideal noiseless operation on a quantum computer can be expressed by a unitary channel $\mathcal{U}_{I}(\rho)=U_{I}\rho U_{I}^{\dagger}$, with $U_{I}$ being some unitary matrix. A simple model for the noisy implementation of this operation on real hardware $\mathcal{E}_{\mathcal{P}}$ consists of the concatenation of a noise modeling Pauli channel $\mathcal{P}$ and the ideal unitary,
$\mathcal{E}_{\mathcal{P}}\left(\rho\right)=\mathcal{P}\!\left(\,\mathcal{U}_{I}(\rho)\right)=\sum_{i}{p(i)P_{i}U_{I}\rho U_{I}^{\dagger}P_{i}},$ (1)
where the Pauli error rates $p(i)$ form a probability distribution over the $n$-qubit Pauli operators $P_{i}\in\mathbb{P}^{\otimes n}=\{I,X,Y,Z\}^{\otimes n}$. Despite the substantial expressibility restrictions of the model by only considering Pauli operators, a wide range of unital incoherent noise, including dephasing and depolarization, can be described by Pauli channels. In this framework, a $n$-qubit noisy operation is described at most by $4^{n}$ Pauli error rates. To reduce this scaling from exponential to polynomial, we follow the widely used and experimentally motivated approach of only regarding locally correlated noise, i.e., considering $n$-qubit Pauli operators $P_{i}$ that act nontrivially on $l$ qubits only [24, 7]. These $l$ qubits are selected with respect to the qubits’ connectivity in the quantum computing architecture. This work focuses on a correlation length $l=2$.
Pauli noise models are appealing for different physical and practical reasons. Depolarization and dephasing are often dominating the qubit systems’ interaction with the environment [34, 35]. Quantum error correction also leads to noise better approximated by Pauli channels [36, 37, 38]. Additionally, Pauli channels are interesting for theoretical work as they can be efficiently simulated classically [39]. Most importantly, however, any noise can be projected onto a Pauli channel by applying Pauli twirling [27, 28, 20, 29, 30, 31]. If $\mathcal{U}_{I}$ is a Clifford operation Pauli twirling can be achieved by stochastically introducing Pauli gates before and after the ideal operation. Enforcing the structure of the model given in Eq. (1) by twirling allows the mitigation of any unknown general noise channel with schemes tailored to Pauli noise, such as PEC. Despite the sampling and circuit overhead due to the probabilistic nature of PEC and twirling [27, 28, 24], this mitigation approach is well-suited for near-term devices because it is implementable with low-depth circuits and relies on Pauli gates [8, 24].
However, in the context of characterization where the goal is to examine the physical noise sources of a system, altering the noise structure using twirling is often not desired, as it obscures the distinction between coherent and incoherent noise. To illustrate the transformative effect of Pauli twirling and motivate the extension of the Pauli noise model in Eq. (1) for characterization purposes, we perform an echo-experiment, shown in Fig. 1 (similar to Ref. [40]), on an IBMq processor. The circuits consist of preparing the state |$\ket{0+}$, an even number of $C\!X$ gates, and measuring the evolved state in the same Pauli basis as it was prepared. As two $C\!X$ gates correspond to the identity, in the noiseless case all measured expectation values are independent of the number of repetitions and equal to $1$. In the exclusive presence of Pauli noise, purely exponential decay of the expectation values, characteristic of incoherent error processes, is expected [20, 21, 22, 28, 41]. This behavior (red dotted curves) can be observed by adding Pauli twirling to the circuits. In the absence of twirling (blue dashed curves), oscillations appear, a feature that exceeds the expressibility of the Pauli noise model in Eq. (1).
Figure 1: Noise examination on identity circuit. The data is obtained on ibm_lagos and plots the expectation value of the three two-qubit Pauli operators $P_{IX},P_{ZI},P_{ZX}$ against the number of noisy identities $m$ implemented by two $C\!X$ gates. In the experiment, we prepare and measure the system in the state |$\ket{0+}$ using one Hadamard gate $H$. The blue dashed curves correspond to the unchanged noise. The dotted red curves are obtained using Pauli twirling scheme that projects the noise to Pauli noise (see Eq. (1)). Each circuit is measured $4096$ times and the twirled results are averaged over $32$ different twirling gates $P$. A readout mitigation scheme [42] is applied in post-processing based on the reported readout error rates of IBM Quantum. The results for different input states and superconducting hardware platforms available on IBMq are similar.
The leading cause for the oscillations in Fig. 1 is coherent noise. We extend the Pauli noise model in Eq. (1) with a unitary channel $\mathcal{U}_{\theta}$ (see Fig. 2) intended to explain the oscillatory behavior observed in Fig. 1
$\mathcal{E}_{\mathcal{P},\theta}\left(\rho\right)=\mathcal{U}_{\theta}\!\left(\,\mathcal{E}_{\mathcal{P}}\!\left(\rho\right)\right)=U_{\theta}\sum_{i}{p(i)P_{i}U_{I}\rho U_{I}^{\dagger}P_{i}U_{\theta}^{\dagger}}$ (2)
with $U_{\theta}$ being a unitary operator parameterized by a Hermitian matrix $H_{\theta}$ as ${U_{\theta}=e^{-iH_{\theta}}}$. This matrix can be expressed in the Pauli basis as $H_{\theta}=\sum_{k}\theta_{k}P_{k}$, with real coefficients $\theta_{k}$ and Pauli matrices $P_{k}$. Assuming small noise, i.e., $U_{\theta}$ is close to identity, we can rewrite the coherent noise as $U_{\theta}\approx\prod_{k}\exp\!\left(-i\theta_{k}P_{k}\right)$, which is a valid approximation up to $\mathcal{O}(\theta^{2})$. Each product term reads
$\exp\!\left(-i\theta_{k}P_{k}\right)=\cos\left({\theta_{k}}\right)\mathbb{I}-i\sin\left(\theta_{k}\right)P_{k},$ (3)
and represents in the Bloch sphere picture a rotation around the $k$-axis by the angle $2\,\theta_{k}$, see Fig. 2. To reach a scalable model, we consider, as for the Pauli noise model in Eq. (1), locally correlated noise operators, namely single- and two-qubit rotations among neighboring qubits according to their connectivity in the hardware. Therefore, the coherently rotated Pauli noise model defined in Eq. (2) consists of nine two-qubit rotations for each neighboring qubit pair and three single-qubit rotations for each qubit, leading to a magnitude of these rotations of $\lvert\theta\rvert=15$.
Note that this model is not universal. Nonunital noise [43], such as decay, surpasses its expressibility. In Ref. [44], the authors analyze the decomposition of general noise into its coherent and incoherent components, investigating how each part contributes to a bound on the total average process fidelity. They show that, for systems with more than one qubit, a physical unitary channel may not fully account for the channel’s rotational components. Pauli noise does not introduce such rotational components, due to their diagonal Pauli-Liouville representation [45]. Consequently, when the system’s incoherent noise is assumed to be Pauli noise, such as dephasing, only the presence of additional coherent noise can lead to oscillatory features in the noise profile. In the discourse regarding the influence of coherent noise on error correcting schemes and the fault tolerance threshold, similar noise models to Eq. (2) are utilized in Refs. [46, 47, 48]. However, those models are limited to specific types of Pauli and coherent noise, and the works do not focus on the characterization of a noisy process but on analyzing the properties of a process with the respective noise model.
Figure 2: Bloch sphere transformations. Image of the Bloch sphere on the left under the action of a Pauli channel $\mathcal{P}$ (middle), and an additional unitary transformation $\mathcal{U}_{\theta}$ (right).
III Characterization
We now present our novel protocol for characterizing all the parameters of the coherently rotated Pauli noise model introduced in Eq. (2). The protocol relies on the assumption that the amplitude of both coherent and incoherent noise is small.
We express the coherent noise characterization protocol in the Pauli transfer matrix (PTM) formalism [49]. As the Pauli matrices build a complete basis of the operator space, a $n$-qubit state $\rho$ can be vectorized $\left|\rho\right)\in\mathbb{R}^{4^{n}}$ in the Pauli basis with the $i$’th element $\left|\rho\right)_{i}$ being their expectation values ${\left|\rho\right)_{i}=\frac{1}{2^{n}}\mathrm{tr}\left({P_{i}\rho}\right)}$. These $4^{n}$ expectation values can be estimated from $3^{n}$ measurements of $\rho$ in different Pauli bases [50, 51, 52, 7, 53, 54]. For a single qubit, this state representation is known as the Bloch vector [55]. In this formalism, a channel $\mathcal{E}$ is represented by a matrix $T\in\mathbb{R}^{4^{n}}\times\mathbb{R}^{4^{n}}$ that acts on the vectorized states according to $\left|\mathcal{E}\!\left(\rho\right)\right)=T\left|\rho\right)$. The matrix $T$ is called the PTM, with elements $T_{ij}=\mathrm{tr}\left(P_{i}\,\mathcal{E}\!\left(P_{j}\right)\right)$. In the context of the coherently rotated Pauli noise model, the diagonal of the PTM matrix encompasses information about the incoherent Pauli noise channel mixed with the second-order terms of the coherent noise. The first-order terms of the coherent noise are contained in the off-diagonal elements without any contribution from Pauli noise.
Before addressing the characterization protocol applicable to a quantum process modeled by the coherently rotated Pauli noise model in Eq. (2), we present the case of estimating $U_{\theta}$ for a 2-qubit process described by $\mathcal{E}_{\theta}\!\left(\rho\right)=U_{\theta}\rho U_{\theta}^{\dagger}$, meaning that we disregard both the Pauli noise $\mathcal{N}$ and the ideal unitary $\mathcal{U}_{I}$ channels. The Pauli transfer matrix $T_{\theta}$ corresponding to $\mathcal{E}_{\theta}$ is a square matrix of order $16$, parameterized by $15$ single- and two-qubit rotation angles. While the first row and column are trivial and independent of $\theta$, in a first-order approximation in $\theta$ (see App. A for a second-order approximation), the other elements of this matrix are given by
$\left(T_{\theta}\right)_{ab}\approx\delta_{ab}-\frac{i}{4}\sum_{k}\theta_{k}\mathrm{tr}\left(\left[P_{a},P_{b}\right]P_{k}\right).$ (4)
By experimentally estimating the vector representation of a set of states $\{\rho_{i}\}$ and their propagation $\{\mathcal{E}_{\theta}\!\left(\rho_{i}\right)\}$ we get an estimate $\theta$ by solving the minimization problem
$\hat{\theta}=\operatorname*{\arg\!\min}_{\phi}\sum_{i}\big{\lVert}\left|\mathcal{E}_{\theta}\!\left(\rho_{i}\right)\right)-T_{\phi}\left|\rho_{i}\right)\big{\rVert}^{2}.$ (5)
Note that compared to full-process tomography on two qubits the number of real parameters is reduced from $240$ to $15$ real parameters, and for a system with $l$-qubits from $\mathcal{O}\left(16^{l}\right)$ to $\mathcal{O}\left(4^{l}\right)$. In App. B we provide the closed-form solution of the minimization in Eq. 5 for a first-order approximation in $\theta$.
The strategy of the characterization protocol for the entire $n$-qubit system is to apply this estimation procedure to every involved two-qubit subsystem according to their connectivity in the hardware. Three points must be considered when moving from the discussed simple case to the model given in Eq. (2): first, the action of the ideal unitary $\mathcal{U}_{I}$; second, the presence of the Pauli noise $\mathcal{N}$; and lastly, the impact of two-qubit noise on qubits that are outside of the considered subsystem.
The effect of the ideal unitary channel $\mathcal{U}_{I}$ can be incorporated in post-processing after estimating the state $\left|\rho\right)$. For an arbitrary $U_{I}$, this procedure requires a full-state tomography of the system [39]. To preserve scalability, we restrict to operations, separable into multiple single- and two-qubit unitaries $U_{I}=U_{1}\otimes U_{2}\otimes...\otimes U_{l}$. This requirement is fulfilled, e.g., in the characterization of so-called gate layers, i.e., $n$-qubit operations consisting of single- and two-qubit gates that can be executed simultaneously in experiments.
[…]
Lastly, when focusing exclusively on two-qubit subsystems, one has to address the correlated multi-qubit noise between a qubit in the regarded subsystem and the external ones. To handle this problem in the estimation of $\left|\rho\right)$ and $\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)$, we prepare the surrounding qubits in different basis states for every single experiment run. This technique isolates the two desired qubits and randomizes the effect of the noise introduced outside the regarded subsystem, similar to twirling schemes.
The protocol for the estimation of coherent noise for a two-qubit system is as follows (for more than 2-qubits see App. D):
  1. 1.
    Choose a positive integers $j$.
  2. 2.
    Initialize the $n$-qubit system in |$\ket{0}^{\otimes n}$.
  3. 3.
    Prepare the system in the eigenstate of a random $n$-qubit Pauli operator $P$ by applying to each qubit $H$, $HX$, $SH$, $SHX$, $X$ or $\mathbb{I}$, obtaining a state |$\ket{\rho}$, with $P\ket{\rho}=\pm\ket{\rho}$.
  4. 4.
    Apply $j$ repetitions of the unitary $U$ to the circuit, resulting in $\mathcal{E}_{\theta}^{j}\!\left(\rho_{j}\right)$.
  5. 5.
    Perform state tomography and analytically reverse the action of the ideal unitary $\mathcal{U}_{I}$. This step results in the state description of $\left|\mathcal{U}_{I}^{\dagger}\,\mathcal{E}_{\theta}^{j}\!\left(\rho_{j}\right)\right)$ and $\left|\mathcal{E}_{\theta}^{j}\!\left(\rho_{j}\right)\right)$.
  6. 6.
    Repeat steps $4$ and $5$ for $j+1$.
  7. 7.
    Repeat steps $1$ to $6$ $L$ times.
  8. 8.
    Aggregate all pairs $\left\{\left|\mathcal{E}_{\theta}^{j}\!\left(\rho_{j}\right)\right),\left|\mathcal{U}^{\dagger}_{I}\,\mathcal{E}_{\theta}^{j+1}\!\left(\rho_{j}\right)\right)\right\}$ for all prepared states and repetitions $j$ and perform the minimization of Eq. (5).
Figure 3: Characterization of the coherent noise introduced by executing a gate layer on the $7$-qubit IBMq Processor ibm_lagos [32]. The gate layer is drawn on top of the connectivity tree (qubit connectivity in the quantum computing architecture), where each big circle marks a different qubit. The gate layer consists of two $C\!X$ gates between qubits $1$ and $2$, and qubits $6$ and $7$, a Hadamard ($H$) gate on qubit 3, a $\sqrt{X}$ ($S\!X$) gate on qubit $4$ and a $X$ gate on qubit $5$. The three pieces inside the qubit representing circles represent the single-qubit coherent errors, and the nine small circles between two qubits display the two-qubit coherent cross-talk errors. All errors are measured in radians and the absolute value is plotted.
We now use the presented protocol to characterize the coherent noise induced by an arbitrary gate layer executed on the IBMq Falcon Processor ibm_lagos. The result is summarized in Fig. 3, where we show the absolute values of the angles of the characterized single- and two-qubit coherent errors for an arbitrary gate layer. The layer is composed of two $C\!X$ gates and three single-qubit gates, namely $H$, $\sqrt{X}$, and $X$. Considering the connectivity of the $7$ qubits in the hardware and our assumption that the noise is only locally correlated, we prepare $216$ initial states, all being product states of |$\ket{0}$, |$\ket{1}$, |+=(|$\ket{+}=(\ket{0}+\ket{1})/\sqrt{2}$, |=(|$\ket{-}=(\ket{0}-\ket{1})/\sqrt{2}$, |$\ket{i}=(\ket{0}+i\ket{1})/\sqrt{2}$ and |$\ket{-i}=(\ket{0}-i\ket{1})/\sqrt{2}$. Each two-qubit subsystem is prepared in $36$ states, and for each of those states, the environment is randomized with $6$ additional states. We execute 00 to $3$ repetitions of the layer, amounting to $4$ circuits per preparation. Finally, bypassing the effect of the unitary gate layer $\mathcal{U}_{I}$ in the post-processing necessitates performing state tomography on up to three-qubit subsystems, leading to $3^{3}$ Pauli measurements per circuit. Overall, we run $328216\cdot 4\cdot 27=23^{\prime}$ 328 different circuits with $128$ shots each, leading to a total running time in the system of about $30$ minutes. A more detailed description of the procedure is given in App. D.
Figure 4: Noise mitigation according to the coherently rotated Pauli noise model in Eq. (2) on ibm_lagos. The plot shows the estimated expectation values of all 9 two-qubit Pauli operators for the initial state $\ket{++}$ after $m$ repetitions of a noisy identity (a) without mitigation, (b) with coherent error mitigation based on data of (a), and (c) with probabilistic error cancellation based on data of (b) and coherent error mitigation. $\ket{++}$ is an eigenstate of the operators $P_{I\!X}$, $P_{X\!I}$, and $P_{X\!X}$, in the noiseless case, their expectation values are $1$. The expectation values of the other 12 Pauli operators (red) are 00. The circuit diagrams contain the preparation of each qubit in one of the states |${\ket{0},\ket{+},\ket{i}}$, by applying Hadamard ($H$) and/or phase ($S$) gates. The evolved states are measured in $9$ bases to estimate all $16$ Pauli expectation values. In (c), specific Pauli operators ($P$) are inserted between the circuit elements according to the PEC method.
The used IBMq platforms report readout errors on the order of 1%, caused by relaxation, imperfect coupling to the readout resonator, and signal amplification errors [56]. To counter the influence of these errors, we apply a post-processing readout mitigation scheme in all executed experiments [42]. While the introduced protocol is not resistant to general state preparation and measurement errors, probabilistic errors that can be modeled by a Pauli channel do not disturb the characterization for the same reasons that the coherent noise characterization is robust against Pauli noise. This is described in detail in App. C. Furthermore, systematic, coherent, preparation errors also don’t affect the protocol, as can be seen in Eq. 5 where the estimation does not depend on information about the intended prepared state.
From Fig. 3, we conclude that the most significant two-qubit coherent errors occur between qubits $1$ and $2$ and qubits $6$ and $7$. Those are associated with the Pauli $P_{Y\!Z}$ error. This result is expected, as those are the pairs of qubits on which the $C\!X$ gates are acting. This Pauli error has already been noted in the analysis of the echoed cross-resonance gate used to realize the $C\!X$ gate [57]. The predominant coherent errors are the single-qubit ones. We note no strong bias towards a specific single-qubit rotation error, and those errors could be related to the hardware implementation of each single-qubit gate.
We conduct the characterization of the same device and gate layer as in Fig. 3 three times, with time differences between the runs being 6 hours and 19 days. The results available in App. E show that the leading coherent errors in the system do not change over time. The most significant single-qubit coherent errors change by no more than $10\%$ and the two-qubit coherent rotation errors by maximally $0.018$ radians. This result suggests that the platform suffers from systematic coherent errors not influenced by the daily re-calibration of the system.
IV Mitigation
In principle, the characterization results in Fig. 3 may originate from overfitting the non-ideal process to the introduced model. Indeed, it is unknown whether the coherently rotated Pauli noise model in Eq. (2) is expressive enough to resemble the main characteristics of the probed IBMq process. To verify that the model captures the dominant noise contributions of a process on the probed device, we examine whether the output of the presented protocol facilitates significant noise mitigation
Returning to the unmitigated circuits in Fig. 1, our goal is to suppress the observed oscillatory behavior in the echo-experiment by correcting the characterized coherent errors at the circuit level. In case of a successful correction, according to the introduced model, one would expect a purely exponential decay of the expectation values with the number of process repetitions due to the Pauli noise. From this exponential decay, we can then characterize the Pauli error rates [7, 24] and mitigate these noise components using PEC [8]. If the final measurement statistics after coherent error mitigation and PEC closely follow the ideal statistics, we demonstrated that the proposed model is physical, and the characterization protocol provides valuable insights.
Any unitary channel is reversible. Therefore, in our model, the effect of the coherent noise can be canceled by single- and two-qubit rotations with opposite rotation angles than the characterized ones. We introduce a parameterized circuit element (see Fig. 7) capable of representing all needed rotations while corresponding to the identity if all parameters are set to zero (see App. F). We use this circuit element to run echo-experiments for three different situations: unmitigated; coherent-error mitigated; and coherent-error mitigated combined with PEC. The results are presented in Fig. 4. Fig. 4(a) and (b) were obtained with 2048 shots per circuit, whereas for Fig. 4(c), an ensemble of 280 altered circuits with 512 shots each was executed per experiment.
For the unmitigated case in Fig. 4(a), we see two main characteristics: Firstly, the expectation values of the 12 Pauli operators that should be ideally 00 are spreading. Secondly, the expectation values that ideally equal $1$, meaning that the states are eigenstates of the corresponding Pauli, decay. Note that the expectation value of the identity in the plot is not based on actual experiments and is always $1$ due to the trace constraints on density matrices.
Applying the presented coherent noise characterization protocol on the data of Fig. 4(a) leads to a first estimate of the rotation angles of the coherent errors $\hat{\theta}_{0}$. To further improve the characterization, we run the characterization scheme with $\vec{\theta}=\hat{\theta}_{0}$ and obtain $\hat{\theta}_{1}$. This iterative approach is expected to improve the characterization because it compensates for violations of the assumed commutation of single- and two-qubit rotation errors. Fig. 4(b) shows the result of the final mitigation iteration with the estimate $\vec{\theta}=\hat{\theta}_{1}$ using the gate-based mitigation circuit from Fig. 7. The observed changes in the angles between the first and the last iteration were about an order of magnitude smaller than the biggest coherent errors and thus supporting the validity of the taken approximation. In Fig. 4(b), the mitigation effect is clearly visible, as the spreading behavior of the expectation values that are ideally 00 is largely suppressed. The expectation values of the Pauli operators, of which the prepared state is a $+1$ eigenstate, are higher in the mitigated case compared to the unmitigated estimation. However, the coherent noise mitigation does not retrieve their ideal-case expectation value of $1$, as according to our model, the statistic is still affected by Pauli noise.
Figure 5: Influence of the coherent error mitigation on fidelity measurement between the ideal state and states evolved by the noisy identity. From the data of Fig. 8, the density matrix of the evolved state is estimated for each number of repetitions of the noisy identity and each of the 9 prepared states. The plot shows the fidelity between these state estimates and the ideally prepared state. The thin lines represent the evolution of the fidelity with the number of repetitions of the noisy identity for each state, and the thick lines are the average of the 9 thin lines.
To show that the examined noise follows the coherently rotated Pauli noise model in Eq. (2), we estimate the Pauli error rates from the mitigated curves of Fig. 4(b) and apply PEC to mitigate the Pauli noise. Note that different from the above presented coherent noise correction, PEC is a mitigation technique that does not correct deterministically individual circuits (as Pauli channels are not reversible) but allows noise-free estimation of expectation values from circuit ensembles. It is important to emphasize that compared to previous works demonstrating PEC [58, 24], we do not use Pauli twirling and, thus, avoid the associated circuit and sampling overhead. The Pauli eigenvalues of the Pauli noise channel are estimated by fitting exponential curves to the expectation values that are supposed to be one in the ideal case. The Pauli error rates can then be found by a Walsh-Hadamard transform [7]. We follow the PEC protocol of van den Berg et al. [24] to mitigate the incoherent noise. It consists of applying the non-physical inverse of a Pauli channel by running an ensemble of unitary processes composing this inverse. Those results are weighted according to the Pauli error rates and combined in post-processing. The results of the echo-experiments with coherent error mitigation and PEC are displayed in Fig. 4(c). Compared to Fig. 4(b), the data are noisier, which we suspect to be due to the PEC sampling overhead.
The obtained measurement statistics clearly resemble the main features of the ideal measurement statistics. Therefore, we can conclude that the introduced noise model explains the leading noise terms in the examined system and that the presented characterization protocol generates valid physical insights into the noise. The data of other input states is shown in App. G. Similar characterization and mitigation results were obtained on ibm_jakarta.
Based on the data from Fig 4(a) and (b), we can reconstruct the density matrices for the $9$ prepared initial states at every repetition of the noisy identity. We then estimate their fidelity with their ideal initial state. The standard state fidelity formula is used $F(\sigma,\rho)=\mathrm{tr}\left(\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}\right)^{2}$. We observe in Fig. 5 that the mitigation decreases the average infidelity by about a factor of two.
V Generalization of the protocol
We motivated our coherently rotated Pauli noise model and devised the protocol for its characterization in a manner agnostic toward the choice of the qubit-based quantum computing architecture. The same procedure can be employed to characterize coherent errors in superconducting devices as we demonstrated in this work, but also in trapped ions, neutral atoms, or any other platforms, as it only requires the preparation of elementary Pauli states without the need for entanglement.
Furthermore, we can readily extend the proposed model to qudit-based quantum processors, which are also susceptible to coherent errors [59, 60, 61, 62, 63]. In these systems, Pauli $X$ and $Z$ operators are generalized as ${X_{d}\ket{s}=\ket{s+1\,\,\mathrm{mod}\,\,d}}$ and ${Z_{d}\ket{s}=\omega^{s}\ket{s}}$, respectively. Here, |$\ket{s}$ represents a qudit computational state with $s\in\{1,2,\ldots,d\}$ and $\omega=\exp(i2\pi/d)$. Consequently, the Pauli noise model for a $n$-qudit system can be expressed similarly to Eq. (1) but with $P_{i}\in\{(X_{d})^{p}(Z_{d})^{q}\,\,|\,\,q,p\in\{1,2,\ldots,d\}\}^{\otimes n}$ representing the generalized Pauli operators. Since the generalized Pauli operators exhibit similar properties to qubit Pauli gates (such as being unitary 1-designs and normalized by the qudit Clifford group), the Pauli transfer matrix formalism remains valid in the qudit framework. It thus enables the same twirling and characterization techniques employed in the qubit case to be utilized [59, 60, 61, 62, 63]. Moreover, this implies that the unitary errors can be effectively modeled by the coherently rotated Pauli noise model described in Eq. (2) and that the characterization protocol presented in Sec. III can be utilized. Additionally, the expansion of the protocol to generalized Pauli operators is also interesting for the characterization of coherent leakage of the qubit. Coherent leakage involves a unitary interaction between the computational subspace and the neglected part of the system’s Hilbert space or its environment [64]. Leakage back to the system introduces noise that is correlated to a past state of the system, introducing memory effects [65]. By expanding the system description to states in the leakage subspace, it is possible to describe the coherent leakage processes equally to coherent noise within the computational subspace. However, this demands precise control or generalized measurements of the leakage subspace, which must be clearly defined and localized.
Similarly, our characterization protocol can be expanded to encompass the logical level, as logical qubits can be susceptible to unitary noise often caused by coherent errors at the physical level [66, 67]. To extend our protocol, it becomes necessary to prepare and measure logical Pauli basis states. These operations are inherent to various error correcting codes, thereby enabling the generalization of our protocol.
VI Conclusion and Outlook
In this paper, we introduced a characterization technique to estimate coherent errors in noisy quantum devices. An existing Pauli noise model is extended to incorporate coherent errors by adding unitary rotations to the model. We devise a procedure to learn the unknown parameters of this coherently rotated Pauli noise model. The scalability of our approach is affirmed by assuming a limited correlation length of the coherent noise. As the results of the characterization of a physical quantum processor showed, the protocol is feasible to be executed on an actual device. For the tested device, the characterization revealed the leading coherent errors introduced by the probed gate layer. Notably, these characterized coherent errors remained stable over multiple re-calibration cycles, suggesting a systematic nature. To verify the validity of our noise model, we mitigated the coherent errors and the characterized Pauli noise. The coherent error mitigation was achieved with a gate-level correction scheme for two-qubit coherent errors. To counter the Pauli noise, we applied probabilistic error cancellation. The combination of the two techniques resulted in almost ideal measurement statistics. This finding shows that the model is sufficiently expressible to explain the main noise contributions of a specific process on an actual processor. Moreover, the successful mitigation of the characterized errors confirms the applicability of the developed protocol. Our protocol is hardware agnostic and thus expected to be executable on a variety of hardware platforms.
Our work offers several promising avenues for future research. The method can enhance mitigation schemes that rely on detailed knowledge of the noise structure. One such approach is optimizing the resilience to coherent noise during circuit transpilation based on the noise insights for each gate or gate layer [68, 69, 70, 71, 72, 71]. A clear example of this is the hidden inverses method [69, 70], which mitigates the impact of coherent errors by compiling circuits in a way that induces destructive interference among the coherent errors of the gates. However, implementing hidden inverses requires knowledge of the physical nature of these errors, which in Ref. [70] was obtained through gate-set tomography. Integrating our approach into the calibration process can provide valuable information to the compiler, enabling it to leverage the hidden inverses technique while actively inverting the non-corrected coherent errors. Similarly, our technique opens the door to studying and developing noise-aware versions of advanced mitigation techniques such as dynamical decoupling or composite pulses. Moreover, information on coherent errors facilitates Pauli conjugation [73]. An efficient noise twirling method that limits sampling overhead compared to common Pauli twirling and, therefore, diminishes the reduction of the threshold of error-correcting codes.
Finally, identifying the nature and strength of the errors is also crucial for the fault-tolerant operation of a quantum processor unit at the logical level [66, 74, 75, 38]. While quantum error correction protocols transform local unitary errors into correctable probabilistic ones [36, 37], coherent noise can lead to significantly larger worst-case logical errors than only incoherent noise [76, 35, 9, 47, 66, 67, 37, 77, 75, 38]. The knowledge gained from coherent errors, which effectively represents soft information, can then be leveraged to enhance the performance of quantum error-correcting codes at both the decoding and encoding stages. Indeed, while optimal decoders are notoriously hard to find [78], recent studies have demonstrated the benefit of soft information in improving existing decoders and overall quantum error correction performance [79, 80, 81]. At the encoding level, it is possible to design codes to be naturally robust against specific coherent noises, as shown in Refs. [82, 83]. Thus, our protocol has the potential to be a diagnostic tool that could help to decide on the next higher-level encoding.
In the long run, the presented characterization protocol paves the way for the development of improved hardware platforms by identifying the limiting imperfections. Yet, already in the near term it can serve as a tool for various applications that rely on knowledge of the existing noise structure, such as quantum error correction or hardware-aware algorithm design.
Acknowledgements.
The authors thank Jonathan Home, Elias Zapusek, Roland Matt, Jeremy Flannery, and Luca Huber for helpful comments and discussions throughout the project. This work was supported by the Swiss National Science Foundation (SNSF) through the National Centre of Competence in Research - Quantum Science and Technology (NCCR QSIT) grant 51NF40–160591. I.R. and F.R. acknowledge financial support by the Swiss National Science Foundation (Ambizione grant no. PZ00P2__\__186040). We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
Appendices
Appendix A Second order approximation in $\theta$ for PTM elements
To find the PTM element of $U_{\theta}$ up to the second order of $\theta$, we first expand $U_{\theta}=\Pi_{k}\text{exp}(-i\theta_{k}P_{k})$ using a second order approximation of Eq. (3):
$U_{\theta}$ $\approx\prod_{k}\left(1-\frac{\theta_{k}^{2}}{2}\right)\mathbb{I}-i\theta_{k}P_{k}$ (6)
$\approx\frac{2-\Sigma_{k}\theta_{k}^{2}}{2}\mathbb{I}-i\sum_{k}\theta_{k}P_{k}-\sum_{k}\sum_{l>k}\theta_{k}\theta_{l}P_{k}P_{l},$
We now use this approximation to evaluate the elements $\left[T_{\theta}\right]_{ab}$ of the PTM up to the second order of $\theta$:
$\left[T_{\theta}\right]_{ab}=\frac{1}{2^{n}}$ $\mathrm{tr}\Bigl{(}P_{a}U_{\theta}^{\dagger}P_{b}U_{\theta}\Bigr{)}$ (7)
$\approx\frac{1}{2^{n}}\mathrm{tr}\Biggl{(}P_{a}$ $P_{b}\left(1-\sum_{k}\theta_{k}^{2}\right)-i\sum_{k}\theta_{k}\left[P_{a},P_{b}\right]P_{k}$
-- $\sum_{k}\sum_{l>k}\theta_{k}\theta_{l}\left(P_{a}P_{l}P_{k}P_{b}+P_{a}P_{b}P_{k}P_{l}\right)$
$+$ $\sum_{k}\sum_{l}\theta_{k}\theta_{l}P_{a}P_{k}P_{b}P_{l}\Biggr{)}$
We can split the last term of the equation into the cases where $k>l$, $k<l$ and $k=l$ and simplify to:
$\sum_{k}\sum_{l}\theta_{k}\theta_{l}P_{a}P_{k}P_{b}P_{l}=\sum_{k}\theta_{k}^{2}P_{a}P_{k}P_{b}P_{k}$ (8)
$+\sum_{k}\sum_{l>k}\theta_{k}\theta_{l}\left(P_{a}P_{k}P_{b}P_{l}+P_{a}P_{l}P_{b}P_{k}\right)$
Using that $P_{a}P_{k}P_{b}P_{k}=(-1)^{\langle k,b\rangle}P_{a}P_{b}$, where $\langle k,b\rangle$ is 00 if $P_{b}$ and $P_{k}$ commute and $1$ otherwise, and the cyclic property of the trace leads to:
$\left[T_{\theta}\right]_{ab}\approx\frac{1}{2^{n}}\mathrm{tr}\Biggl{[}$ $\left(1\!-\!\sum_{k}\theta_{k}^{2}\left(1\!-\!(-1)^{\langle b,k\rangle}\right)\right)P_{a}P_{b}$ (9)
$-i\sum_{k}\theta_{k}\left[P_{a},P_{b}\right]P_{k}$
$+\sum_{k}\sum_{l>k}\theta_{k}\theta_{l}\left[P_{a},P_{l}\right]\left[P_{b},P_{k}\right]\Biggr{]},$
where all terms of order in $\theta$ larger than two are dropped. Finally, we reach the approximation of the PTM element of $U_{\theta}$ up to the second order in $\theta$:
$\left[T_{\theta}\right]_{ab}\approx$ $\left(1-\sum_{k}2\theta_{k}^{2}\langle a,k\rangle\right)\delta_{ab}$ (10)
$-\frac{i}{2^{n}}\sum_{k}\theta_{k}\mathrm{tr}\left(\left[P_{a},P_{b}\right]P_{k}\right)$
$+\frac{1}{2^{n}}\sum_{k}\sum_{l>k}\theta_{k}\theta_{l}\mathrm{tr}\left(\left[P_{a},P_{l}\right]\left[P_{b},P_{k}\right]\right).$
Assuming the coherent noise is sufficiently weak, we can simplify this expression to obtain the approximation in Eq. (4).
Appendix B Closed form expression for estimation
The right hand side of Eq. (4) can be written equivalently as:
$\left(T_{\theta}\right)_{ij}\approx\delta_{ij}-\frac{i}{4}\theta_{k}\mathrm{tr}\left(\left[P_{i},P_{j}\right]P_{k}\right)=\delta_{ij}+\theta_{k}\,C_{kij},$ (11)
where we use Einstein’s notation for the summation over $k$ and $C$ is a tensor with coefficients $k,i$ and $j$. Applying the channel to an arbitrary state $\left|\rho\right)$ leads to the equation ${\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)_{i}=\left|\rho\right)_{i}+\theta_{k}\,C_{kij}\left|\rho\right)_{j}}$, again applying Einstein’s notation to sum over $k$ and $j$. Defining the matrix $(B_{\left|\rho\right)})_{ik}=C_{kij}\left|\rho\right)_{j}$, the estimation problem boils down to a linear inverse problem of form ${B_{\left|\rho\right)}\,\theta=\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)-\left|\rho\right)}$ where the vectors $\left|\rho\right)$ and $\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)$ can be experimentally estimated and $B_{\left|\rho\right)}$ is a function of $\left|\rho\right)$. However, as $B_{\left|\rho\right)}$ is singular and cannot be inverted. To overcome this we must estimate $B_{\left|\rho\right)}$, $\left|\rho\right)$ and $\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)$ for multiple states $\rho$. Doing this for $N$ states, we formulate the overdetermined system $\tilde{B}\theta=y$, where $\tilde{B}$ of size $15N\times 15$ is the concatenation of all matrices $B_{\left|\rho_{i}\right)}$ and $y$ the concatenation of all vectors $\left|\mathcal{E}_{\theta}\!\left(\rho_{i}\right)\right)-\left|\rho_{i}\right)$. Finally, we estimate $\theta$ using the least square method,
$\hat{\theta}=\left(\tilde{B}^{T}\tilde{B}\right)^{-1}\tilde{B}^{T}y\,.$ (12)
Figure 6: Coherent noise drift. We repeated the characterization result of Fig. 3 at $3$ different times. Subplot (a) shows the baseline data of Fig. 3. Subplot (b) displays the observed changes in the coherent error characterization over a time span of 19 days and subplot (c) exhibits the differences over (6) hours. The experiments were run on the $7$-qubit IBM Quantum Processor ibm_lagos [32] on September 22, 2022, and October 11, 2022.
Appendix C Addressing Pauli noise neglect during coherent noise characterization
This section complements the intuitive reasoning of Sec. III, on why we can neglect the influence of the Pauli noise channel when characterizing the present coherent noise, with a mathematically more rigorous explanation.
The action of a noisy channel modeled by $\mathcal{E}_{\mathcal{P},\theta}(\rho)$ as defined in Eq. (2) can be represented in terms of the vectorized form of the density matrix of $\rho$ and the PTM of the channel. It reads ${\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)=T_{\mathcal{E}_{\mathcal{P},\theta}}\left|\rho\right)=T_{U_{\theta}}T_{\mathcal{P}}T_{U_{I}}\left|\rho\right)}$ where $T_{U_{\theta}},\,T_{\mathcal{P}}\text{ and }T_{U_{I}}$ are the transfer matrices of the coherent, Pauli and the ideal unitary channels, respectively. As stated in the main text, we discuss the case of $T_{U_{I}}$ being the identity. We moreover assume that $\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)$ and $\left|\rho\right)$ can be experimentally estimated. The estimation of the 15 parameters of $T_{U_{\theta}}$ can be written as a minimization problem over the estimator $\hat{T}_{U_{\theta}}$.
Statement 1.
With the assumption $T_{\mathcal{P}}$ and $T_{U_{\theta}}$ being close to the identity. The minimization of the function
$\left\lVert\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)-\hat{T}_{U_{\theta}}T_{\mathcal{P}}\left|\rho\right)\right\rVert^{2}$ (13)
over $\hat{T}_{U_{\theta}}$ is equivalent to the minimization of
$\left\lVert\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)-\hat{T}_{U_{\theta}}\left|\rho\right)\right\rVert^{2}\,\,.$ (14)
Proof.
A noisy identity process described by the coherent noise model, $\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)$ equals $T_{U_{\theta}}T_{\mathcal{P}}\left|\rho\right)$. We show that $\theta^{\prime}=\theta$ minimizes the expression $\lVert T_{\mathcal{P}}T_{U_{\theta}}\left|\rho\right)-T_{U_{\theta^{\prime}}}\left|\rho\right)\rVert$ independent of the choice of $\rho$.
$\min_{\theta^{\prime}}{\left\lVert\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)-{T}_{U_{\theta^{\prime}}}\left|\rho\right)\right\rVert^{2}}$ (15)
$=\min_{\theta^{\prime}}{\left\lVert T_{U_{\theta}}T_{\mathcal{P}}\left|\rho\right)-{T}_{U_{\theta^{\prime}}}\left|\rho\right)\right\rVert^{2}}$
$=\min_{\theta^{\prime}}{{\left|\rho\right)}^{\top}\!\left(T_{\mathcal{P}}^{\top}T_{U_{\theta}}^{\top}-{T}_{U_{\theta^{\prime}}}^{\top}\right)\!\Big{(}T_{U_{\theta}}T_{\mathcal{P}}-{T}_{U_{\theta^{\prime}}}\Big{)}\left|\rho\right)}$
$=\min_{\theta^{\prime}}{\left(\rho\right|}\left(T_{\mathcal{P}}^{\top}T_{U_{\theta}}^{\top}T_{U_{\theta}}T_{\mathcal{P}}-T_{U_{\theta^{\prime}}}^{\top}T_{U_{\theta}}T_{\mathcal{P}}-\right.$
$\left.\qquad\qquad\quad-T_{\mathcal{P}}T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}+{T}_{U_{\theta^{\prime}}}^{\top}{T}_{U_{\theta^{\prime}}}\right)\left|\rho\right)\,.$
Given that the noise channel is close to the identity, we take the approximation that the diagonal matrix $T_{\mathcal{P}}$ commutes with the matrices $T_{U_{\theta}}$ and $T_{U_{\theta}^{\prime}}$. Furthermore, because of because of the unitarity of the coherent rotation channel, $T_{U_{\theta}}^{\top}{T_{U_{\theta}}}=\mathbb{I}$ is valid. Then, the minimization problem boils down to
$\min_{\theta^{\prime}}\left(\rho\right|\left(\mathbb{I}\!+\!T_{\mathcal{P}}^{2}\!-\!T_{\mathcal{P}}\!\left(T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}\!+\!\left(T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}\right)^{\!\!\top}\right)\right)\left|\rho\right)\,.$ (16)
The off-diagonal part of $T_{U_{\theta}}^{\top}T_{U_{\theta^{\prime}}}$ is anti-symmetric. Consequently, $T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}+\left(T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}\right)^{\!\!\top}$ is diagonal and smaller or equal than ${2~{}\mathbb{I}}$. Thereby, all involved matrices are diagonal. From ${1+T_{\mathcal{P}}^{2}\geq 2~{}T_{\mathcal{P}}}$, it follows that the expression in Eq. (16) is minimal when ${T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}+\left(T_{U_{\theta}}^{\top}{T}_{U_{\theta^{\prime}}}\right)^{\!\!\top}=2~{}\mathbb{I}}$. This equation has a unique solution corresponding to $T_{U_{\theta}}={T}_{U_{\theta^{\prime}}}$, which finally implies $\theta=\theta^{\prime}$.
This statement implies that we can optimize for the coherent noise term without considering the present Pauli noise, allowing to apply the closed form solution in Eq. (12) or conduct a least squares optimization method for $\lVert\left|\mathcal{E}_{\mathcal{P},\theta}(\rho)\right)-\hat{T}_{U_{\theta}}\left|\rho\right)\rVert^{2}$.
[…]
Appendix D step-by-step circuit generation
We aim to provide a step-by-step overview of the circuit construction for gate-layer characterization. Each circuit has three main parts: state preparation, the gate layer of interest, and measurement. Since we prepare and measure the states in a Pauli basis, those two parts only require the local gates $X$, $H$, $S$, and $S^{\dagger}$. We choose a set of $L$ states to prepare, ensuring that there is no correlation in the preparation of nearby qubits. These states can be selected randomly; however, as described in Sec. III, we choose a more structured set of $216$ initial states that are all combinations of the states $\ket{+}$, $\ket{-}$, |$\ket{i}$, |$\ket{-i}$, |$\ket{1}$, and |$\ket{0}$ prepared by applying the gates $H$, $HX$, $SH$, $SHX$, $X$, respectively $\mathbb{I}$ to the state |$\ket{0}$. Next, we apply $j$ and $j+1$ iterations of the gate-layer, where $j$ is an integer. At the final measurement stage, we need a basis that allows us to reverse the action of the ideal unitary on each subsystem of connected qubits. For every neighboring pair of qubits, $q_{i}$ and $q_{j}$, the required measurement basis must encompass the subsystem that includes these two qubits along with all other qubits sharing a two-qubit gate with them. To achieve this, we use a complete Pauli measurement basis for each subsystem, which consists of all possible combinations of measuring each qubit in the $x$, $y$, or $z$ basis.
The general procedure for arbitrary system sizes and $k$-local noise is as follows:
  1. 1.
    Choose a positive integer $j$.
  2. 2.
    Initialize the $n$-qubit system in |$\ket{0}^{\otimes n}$.
  3. 3.
    Prepare the system in the eigenstate of a random $n$-qubit Pauli operator $P$ by applying to each qubit $H$, $HX$, $SH$, $SHX$, $X$ or $\mathbb{I}$, obtaining a state |$\ket{\rho}$, with $P\ket{\rho}=\pm\ket{\rho}$.
  4. 4.
    Apply $j$ repetitions of the unitary $U$ to the circuit, resulting in $\mathcal{E}_{\theta}^{j}\!\left(\rho_{j}\right)$
  5. 5.
    Perform state tomography on each relevant $l$-qubit subsystem. With $k$ being the assumed locality of the considered noise, $l$ must be large enough to analytically reverse the action of the ideal unitary $\mathcal{U}_{I}$ for each $k$-qubit subsystem. This step results in the state description of $\mathcal{U}_{I}^{\dagger}\,\mathcal{E}_{\theta}^{j}\!\left(\rho_{j}\right)$ and $\mathcal{E}_{\theta}^{j}\!\left(\rho_{j}\right)$ for each $k$-qubit subsystem.
  6. 6.
    Repeat steps $4$ and $5$ for $j+1$.
  7. 7.
    Repeat steps $1$ to $6$ $L$ times
  8. 8.
    For each $k$-qubit subsystem, aggregate all pairs $\left\{\left|\mathcal{E}_{\theta}^{j}\!\left(\rho_{j}\right)\right),\left|\mathcal{U}^{\dagger}_{I}\,\mathcal{E}_{\theta}^{j+1}\!\left(\rho_{j}\right)\right)\right\}$ for all prepared states and repetitions $j$. Perform the minimization of Eq. (5).
Figure 7: Coherent noise mitigating circuit element. The circuit element comprises 4 $C\!X$ gates, 4 fixed $z$-rotations by $\pi/2$, respectively $-\pi/2$, and 15 parameterized single-qubit rotations. The first 9 rotations (blue) account for different two-qubit coherent errors, and the final 6 rotations (red) mitigate single-qubit errors. Without the first $C\!X$ gate, the circuit element represents a corrected $C\!X$ gate. We represent the element by the symbol on the right-hand side, where $\vec{\theta}$ corresponds to the angles of the 15 parameterized rotations.
Appendix E Coherent noise drift
We conducted the characterization of the gate layer represented in Fig. 3 again after 6 hours as well as 19 days. The shifts in the characterization results between those three experiments are shown in Fig. 6. First, we observe that the leading coherent errors in the system did not change over the 19 days. However, as expected, the changes observed for the two experiments within 6 hours are significantly smaller than the differences over the complete time span. After 6 hours, the maximum deviation for single-qubit coherent errors was $0.009$ radians and $0.004$ radians on average. The two-qubit errors changed maximally by $0.008$ radians and on average by $0.003$ radians. In the total period (i.e. after 19 days), the most significant single-qubit coherent errors have changed by maximally $10\%$ (corresponding to about $0.03$ radians). The two-qubit coherent rotation errors have changed by maximally $0.018$ radians. The average deviation was $0.007$ radians for the single-qubit errors and $0.003$ radians for the two-qubit errors. These results suggest that the platform is affected by systematic coherent errors which are not influenced by the hourly and daily re-calibrations of the system during which qubits’ frequency and readout angle, as well as pulses’ amplitude and phase of basic single- and two-qubit gates, are calibrated [32].
Figure 8: Complete data set from coherent error mitigation. Collected data of the coherent error mitigated experiment of Fig. 3. The blue curves display the unmitigated results as shown in Fig. 3 (a). Red shows the curves of the experiment with coherent mitigation according to Fig. 3 (b). Each of the 9 different panels exhibits a different initial 2-qubit state.
Appendix F Coherent noise correcting circuit element
To conduct echo-experiments that allow for correcting coherent errors we require a parameterized circuit element capable of representing the characterized rotations. For a starting point of the experiment, when all coherent noise correction terms are set to zero, the circuit must correspond to the identity circuit. Single-qubit rotations are readily available in most of the leading quantum computing architectures. To invert the two-qubit rotations, an entangling gate is necessary. We use the $C\!X$ gate as it is a native gate of the chosen platform. By interleaving 15 single-qubit rotations into a structure of four $C\!X$ gates and four fixed single-qubit rotations, as shown in Fig. 7, a circuit element with 15 parameters is created. Each of the 15 parameters, which we will denote as $\vec{\theta}$ corresponds to a different single- or two-qubit rotation. The element corresponds to the identity if all parameters are set to zero.
Appendix G Complementary mitigation data
Fig. 8 displays the complete data set used to create Fig. 5. The influence on the expectation values that ideally are 00 for all number of repetitions is clearly visible for all states, as they disperse much less in the mitigated case. Furthermore, the expectation values of the operators, of which the prepared state is an eigenstate, are closer to the ideal value of $1$ in the mitigated case. While the mitigation did not work equally well for all 9 states, the whole data set gives a clear indication of the positive effect of the presented coherent noise mitigation scheme.